Given: x = tan4t and y = sec4t. Find y’.


1
Expert's answer
2022-02-23T08:27:03-0500
"y'_x=\dfrac{y'_t}{x'_t}"

"x'_t=4\big(\dfrac{1}{\cos^2(4t)}\big)"

"y'_t=4\big(-\dfrac{1}{\cos^2(4t)}\big)(-\sin(4t)"

"y'_x=\dfrac{y'_t}{x'_t}=\dfrac{\dfrac{4\sin (4t)}{\cos^2(4t)}}{\dfrac{4}{\cos^2(4t)}}=\sin (4t)"

"y'_x=\sin(4t)"